Jason La Rue can shampoo the carpet on the main floor of the Sheraton Hotel in 3 hours. Tom Lockheart can shampoo the same carpet in 6 hours. If they work together, how long will it take them to shampoo the carpet?
Answer by nerdybill(7384)

Jason La Rue can shampoo the carpet on the main floor of the Sheraton Hotel in 3 hours. Tom Lockheart can shampoo the same carpet in 6 hours. If they work together, how long will it take them to shampoo the carpet?
Let x = time (hours) it takes for both
then
x(1/3 + 1/6) = 1
multiplying both sides by 6:
x(2 + 1) = 6
x(3) = 6
x = 6/3
x = 2 hours